OpenStudy (anonymous):

Limits! Evaluate using the squeeze theorem. lim x^2 cos 1/x x->0 Help!! I have a lot of these to do so details would be uber helpful

OpenStudy (turingtest):

\(\cos\theta\) is always between -1 and 1, regardless of what \(\theta\) is. This means that\[x^2\le |x^2\cos\theta|\]

OpenStudy (turingtest):

sorry, I meant that\[|x^2\cos\theta|\le x^2\]

OpenStudy (turingtest):

therefore we can write\[-x^2\le x^2\cos(\frac1x)\le x^2\]